Jane wants to purchase a bond with a face value of $8,595 and a bond rate of 7% per year, payable quarterly. The bond has a remaining life of 4 years. If Jane wants to earn at least 10% per year compounded quarterly, what is the maximum price she would be willing to pay to purchase the bond?
Answer #1

face value = 8595

coupon payable = 7% / 4 * 8595 = 150.41

t = 4 * 4 = 16 quarters

I = 10% / 4 = 2.5% per quarter

maximum price of bond = 150.41 * (P/A,2.5%,16) + 8595 * (P/F,2.5%,16)

= 150.41 * ((1 + 0.025)^16-1)/(0.025 * (1 + 0.025)^16) + 8595 * (1 + 0.025)^-16

= 150.41 * ((1.025)^16-1)/(0.025 * (1.025)^16) + 8595 * (1.025)^-16

= 150.41 *13.055003 + 8595 *0.673625

= 7753.41